IN-SITU PHOTOLYSIS OF METHYL IODIDE IN SOLID PARA-HYDROGEN AND SOLID ORTHO-DEUTERIUM

\underline{YUKI MIYAMOTO}, Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan; MIZUHO FUSHITANI, Institute fur Experimentalphysik, Freie Universitat Berlin, Arnimallee 14, 14195, Berlin, Germany; HIROMICHI HOSHINA, Terahertz Sensing and Imaging Laboratory, RIKEN; 519-1399 Aramaki Aoba, Sendai 980-0845, Japan; and TAKAMASA MO-MOSE, Department of Chemistry, The University of British Columbia, Vancouver, B.C., V6T1Z1, Canada.

Products of in-situ photolysis of methyl iodide isolated in solid para-hydrogen and solid ortho-deuterium were investigated by high-resolution FTIR spectroscopy. The ν_{3} and ν_{4} vibrational transitions of a methyl radical and the ^{2}P_{1/2} ←^{2} P_{3/2} transition of an iodine atom were clearly observed in both solids after the photolysis, which indicates a small cage effect in hydrogen crystals. The vibrational transitions of the methyl radical showed a rotational structure subjected to crystal field splittings. Temperature dependence of the spectra allowed us to analyze the spectral fine structure completely based on our crystal field theory.^{a} On the other hand, the magnetic dipole transition of the iodine atom showed a rotational satellite induced by hydrogen molecules. The rotational satellite indictates a strong electron-roton coupling in solid hydrogen. Details of the analysis will be discussed.

aT. Momose, H. Hoshina, M. Fushitani, and H. Katsuki, Vib. Spectrosco. 34, 95 (2004).